# AC-PKAN: Attention-Enhanced and Chebyshev Polynomial-Based Physics-Informed Kolmogorov–Arnold Networks

### v1

@@ Appendix F Impact Statement
Appendix G Supplementary Experiment
G.1 Complex Function Fitting
We evaluated our AC-PKAN Simplified model—which employs only the internal architecture—against PINN (MLP), KAN, and various KAN variants on a complex function interpolation task. Detailed experimental setups and results are provided in Appendices H and I.
As shown in Figure 3, the AC-PKAN Simplified model converges more rapidly than MLPs, KAN, and most KAN variants, achieving lower final losses. While Cheby2KAN and FourierKAN demonstrate faster convergence, our model produces smoother fitted curves and exhibits greater robustness to noise, effectively preventing overfitting in regions with high-frequency variations.Performance metrics are presented in Table 6.
Model rMAE rMSE Loss
Cheby1KAN 0.0179 0.0329 0.0068
Cheby2KAN 0.0189 0.0313 0.0079
MLP 0.0627 0.1250 0.1410
AC-PKAN_s 0.0177 0.0311 0.0081
KAN 0.0145 0.0278 0.0114
rKAN 0.0458 0.0783 0.1867
fKAN 0.0858 0.1427 0.1722
FastKAN 0.0730 0.1341 0.1399
FourierKAN 0.0211 0.0353 0.0063
Table 6: Comparison of test rMAE, rMSE, and training Loss among Models in Complex Function Fitting Experiment
Figure 3: Convergence Comparison of Nine Different Models in Complex Function Fitting Experiment
G.2 Integration with Other External Learning Strategies for Enhanced Performance of AC-PKAN
Model rMAE rRMSE
AC-PKAN + NTK 0.0009 0.0009
PINNs + NTK 0.1397 0.1489
PINNsFormer + NTK 0.0453 0.0484
Appendix H Experiment Setup Details
@@ H.1 Running Time
Model First 5 PDEs Last 3 PDEs
1D-Wave 1D-Reaction 2D NS Cylinder 1D-Conv.-Diff.-Reac. 2D Lid-driven Cavity Heterogeneous Complex 3D
Problem Geometry Point-Cloud
PINN 00:21:14 00:09:07 00:15:20 00:15:12 00:06:39
00:23:30 00:01:08 00:49:31
PINNsFormer 00:44:21 00:04:09 00:58:54 02:06:37
14:01:55 00:13:31
QRes 01:41:34 00:02:10 00:24:39 00:25:46 00:13:04
00:20:50 00:01:46 01:32:24
FLS 01:38:01 00:01:29 00:11:51 00:50:26 00:35:48
00:13:38 00:01:08 03:04:41
Cheby1KAN 03:32:10 00:12:08 04:24:59 01:45:37 00:45:20
00:50:45 00:03:21 02:27:27
Cheby2KAN 05:03:18 01:06:54 05:41:42 03:01:40 00:45:15
01:35:40 00:03:27 05:26:42
AC-PKAN 01:13:01 00:15:16 02:21:40 02:01:59 00:51:47
01:13:11 00:01:04 04:54:24
KINN 25:00:20 03:04:19 14:31:42 02:41:49
01:51:44 00:14:07
rKAN 12:44:16 01:21:25 05:19:04 02:06:36
06:21:00 00:16:06 07:53:25
FastKAN 09:35:51 05:51:21 02:04:42 03:22:39
03:37:57 00:17:23
fKAN 08:20:34 00:13:09 03:01:41 01:54:22 00:47:41
00:52:05 00:06:22 04:04:48
FourierKAN 03:33:46 01:21:50 02:48:50 02:08:08
07:40:43 00:18:26 13:36:48
Table 8: Running times (hh:mm:ss) for all eight PDE experiments. First row: Five simpler PDEs; second row: Three more complex cases.
H.2 Complex Function Fitting Experiment Setup Details
@@
H.3 Failure Modes in PINNs Experiment Setup Details
@@
We set the parameters $\lambda_{1}=1$ and $\lambda_{2}=0.01$. Since the system lacks an explicit analytical solution, we utilize the simulated solution provided in [50]. We focus on the prototypical problem of incompressible flow past a circular cylinder, a scenario known to exhibit rich dynamic behavior and transitions across different regimes of the Reynolds number, defined as $\text{Re}=\frac{u_{\infty}D}{\nu}$. By assuming a dimensionless free-stream velocity $u_{\infty}=1$, a cylinder diameter $D=1$, and a kinematic viscosity $\nu=0.01$, the system exhibits a periodic steady-state behavior characterized by an asymmetric vortex shedding pattern in the cylinder wake, commonly known as the Kármán vortex street. All experimental settings are the same as in [50]. For more comprehensive details about this problem, please refer to that work.
@@
$X$ (Spatial Coordinates) (5000, 2)50002(5000,\,2)( 5000 , 2 ) Contains 5,000 spatial points, each with 2 coordinate values ($x$ and $y$).
$t$ (Time Data) (200, 1)2001(200,\,1)( 200 , 1 ) Contains 200 time steps, each corresponding to a scalar value.
$U$ (Velocity Field) (5000, 2, 200)50002200(5000,\,2,\,200)( 5000 , 2 , 200 ) Contains 5,000 spatial points, 2 velocity components ($u$ and $v$), and 200 time steps. The velocity data of each point is a function of time.
$P$ (Pressure Field) (5000, 200)5000200(5000,\,200)( 5000 , 200 ) Contains pressure data for 5,000 spatial points and 200 time steps.
@@
$\begin{gathered}\frac{\partial u}{\partial t}+\beta\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}-\rho u(1-u)=0,\quad\forall x\in[0,2\pi],\;t\in[0,1]\\ \texttt{IC:}\quad u(x,0)=\exp\left(-\frac{(x-\pi)^{2}}{2(\pi/4)^{2}}\right),\qquad\texttt{BC:}\quad u(0,t)=u(2\pi,t)\end{gathered}$ (128)
In this equation, $\beta$ represents the convection coefficient, $\nu$ is the diffusivity, and $\rho$ is the reaction coefficient. Specifically, we set $\beta=1$, $\nu=3$, and $\rho=5$. The reaction term adopts the well-known Fisher’s form $\rho u(1-u)$, as utilized in [33]. This formulation captures the combined effects of transport, spreading, and reaction dynamics on the state variable $u$.
@@
$x$ (Spatial Coordinates) (10,201, 1)102011(10,201,\,1)( 10 , 201 , 1 ) Represents spatial points uniformly distributed over the domain $[0,2\pi]$.
$t$ (Time Data) (10,201, 1)102011(10,201,\,1)( 10 , 201 , 1 ) Denotes temporal data spanning the domain $[0,1]$ for solution evolution.
$u$ (Solution Values) (10,201, 1)102011(10,201,\,1)( 10 , 201 , 1 ) Contains the computed values of the solution function $u(x,t)$ at corresponding spatial and temporal points.
@@
Out of the total $10,\!201$ data points, the dataset was partitioned into training and test sets. The training data includes boundary points (where $x=0$ or $x=2\pi$) and a random sample of $3,\!000$ interior points, which were used to compute the loss function during model training. The test data consists of the entire remaining dataset, ensuring comprehensive evaluation of the model’s performance.
@@
$\begin{gathered}\mathbf{u}\cdot\nabla\mathbf{u}+\nabla p-\frac{1}{\text{Re}}\Delta\mathbf{u}=0,\quad\forall\mathbf{x}\in\Omega,\;t\in[0,T]\\ \nabla\cdot\mathbf{u}=0,\quad\forall\mathbf{x}\in\Omega,\;t\in[0,T]\\ \texttt{IC:}\quad\mathbf{u}(\mathbf{x},0)=\mathbf{0}\\ \texttt{BC:}\quad\mathbf{u}=(4x(1-x),0),\quad\mathbf{x}\in\Gamma_{1}\\ \mathbf{u}=(0,0),\quad\mathbf{x}\in\Gamma_{2}\\ p=0,\quad\mathbf{x}=(0,0)\end{gathered}$ (129)
@@
$X$ (Spatial Coordinates) (10,201, 2)102012(10,\!201,\,2)( 10 , 201 , 2 ) Contains 10,201 spatial nodes with $(x,y)$ coordinates spanning the cavity domain.
$U$ (Velocity Field) (10,201, 2)102012(10,\!201,\,2)( 10 , 201 , 2 ) Horizontal ($u$) and vertical ($v$) velocity components at $\mathrm{Re}=100$, with no-slip boundary conditions and a moving lid ($y=1$) driving the flow.
$P$ (Pressure Field) (10,201, 1)102011(10,\!201,\,1)( 10 , 201 , 1 ) Pressure values normalized with respect to the reference boundary condition.
@@
The training set comprises 3,000 randomly sampled spatial points with associated velocity and pressure values, while the test set evaluates the model on the full dataset of 10,201 nodes. Boundary conditions are explicitly enforced for the moving lid ($u=4x(1-x),v=0$) and stationary walls ($u=v=0$), with the pressure field satisfying the incompressibility constraint.
@@ 2 $\rightarrow$ 512 (Linear, SinAct)
H.4 PDEs in Complex Engineering Environments Setup Details
@@
$\begin{cases}a_{1}\Delta u(\boldsymbol{x})=16r^{2}&\text{for }r<r_{0},\\ a_{2}\Delta u(\boldsymbol{x})=16r^{2}&\text{for }r\geq r_{0},\\ u(\boldsymbol{x})=\frac{r^{4}}{a_{2}}+r_{0}^{4}\left(\frac{1}{a_{1}}-\frac{1}{a_{2}}\right)&\text{on }\partial\Omega,\end{cases}$ (130)
where $r=\|\boldsymbol{x}\|_{2}$ is the distance from the origin, $a_{1}=\frac{1}{15}$ and $a_{2}=1$ are the material coefficients, $r_{0}=0.5$ defines the interface between the two materials, and $\partial\Omega$ represents the boundary of the square domain $\Omega=[-1,1]^{2}$. The boundary condition is a pure Dirichlet condition applied uniformly on all four edges of the square.
@@
$-\Delta u=0\quad\text{in }\Omega=\Omega_{\text{rec}}\setminus\bigcup_{i=1}^{4}R_{i},$ (131)
@@
$R_{1}$ =$=\left\{(x,y):(x-0.3)^{2}+(y-0.3)^{2}\leq 0.1^{2}\right\},$
$R_{2}$ =$=\left\{(x,y):(x+0.3)^{2}+(y-0.3)^{2}\leq 0.1^{2}\right\},$
$R_{3}$ =$=\left\{(x,y):(x-0.3)^{2}+(y+0.3)^{2}\leq 0.1^{2}\right\},$
$R_{4}$ =$=\left\{(x,y):(x+0.3)^{2}+(y+0.3)^{2}\leq 0.1^{2}\right\}.$
@@
$u$ $=0\quad\text{on }\partial R_{i},\quad\forall i=1,2,3,4,$ (132)
$u$ $=1\quad\text{on }\partial\Omega_{\text{rec}}.$ (133)
@@
$-\mu(z)\Delta u(\mathbf{x})+k(z)^{2}u(\mathbf{x})=f(\mathbf{x})\quad\text{in }\Omega=[0,1]^{3}\setminus\bigcup_{i=1}^{4}\mathcal{C}_{i},$ (134)
@@
AC-PKAN
Linear Embedding: in=2, out=32
ChebyKAN Layers: 4 layers, degree=8
LN Layers: 4 layers, features=64
Output Layer: in=64, out=1
Activation: WaveAct
152,357
QRes
Input Layer: in=2, out=128
Hidden Layers: 5 QRes blocks, units=128
Output Layer: in=128, out=1
Activation: Sigmoid
166,017
PINN
Layer 1: 2 $\rightarrow$ 256, Activation=Tanh
Layer 2: 256 $\rightarrow$ 512, Activation=Tanh
Layer 3: 512 $\rightarrow$ 128, Activation=Tanh
Layer 4: 128 $\rightarrow$ 1
198,145
PINNsformer
d_out=1
d_hidden=128
d_model=8
N=1
heads=2
158,721
FLS
Layer 1: 2 $\rightarrow$ 256, Activation=SinAct
Layer 2: 256 $\rightarrow$ 256, Activation=Tanh
Layer 3: 256 $\rightarrow$ 256, Activation=Tanh
Layer 4: 256 $\rightarrow$ 1
132,609
Cheby1KAN
Layer 1: 2 $\rightarrow$ 32, Degree=8
Layer 2: 32 $\rightarrow$ 128, Degree=8
Layer 3: 128 $\rightarrow$ 64, Degree=8
Layer 4: 64 $\rightarrow$ 32, Degree=8
Layer 5: 32 $\rightarrow$ 1, Degree=8
129,888
Cheby2KAN
Layer 1: 2 $\rightarrow$ 32, Degree=8
Layer 2: 32 $\rightarrow$ 128, Degree=8
Layer 3: 128 $\rightarrow$ 64, Degree=8
Layer 4: 64 $\rightarrow$ 32, Degree=8
Layer 5: 32 $\rightarrow$ 1, Degree=8
129,888
KAN*
Layers: 2 $\times$ KANLinear
Neurons: 9000
Activation: SiLU
60,000*
rKAN
Layer 1: 2 $\rightarrow$ 256, Activation=JacobiRKAN(3)
Layer 2: 256 $\rightarrow$ 256, Activation=PadeRKAN[2/6]
Layer 3: 256 $\rightarrow$ 256, Activation=JacobiRKAN(6)
Layer 4: 256 $\rightarrow$ 128, Activation=PadeRKAN[2/6]
Layer 5: 128 $\rightarrow$ 1
165,411
FastKAN*
FastKANLayer 1: RBF, SplineLinear 16 $\rightarrow$ 2600, Base Linear 2 $\rightarrow$ 2600
FastKANLayer 2: RBF, SplineLinear 20800 $\rightarrow$ 1, Base Linear 2600 $\rightarrow$ 1
75,418*
fKAN
Layer 1: 2 $\rightarrow$ 256, Activation=fJNB(3)
Layer 2: 256 $\rightarrow$ 512, Activation=fJNB(6)
Layer 3: 512 $\rightarrow$ 512, Activation=fJNB(3)
Layer 4: 512 $\rightarrow$ 128, Activation=fJNB(6)
Layer 5: 128 $\rightarrow$ 1
132,618
FourierKAN
Layer 1: 2 $\rightarrow$ 32
Layer 2: 32 $\rightarrow$ 64
Layer 3: 64 $\rightarrow$ 64
Layer 4: 64 $\rightarrow$ 64
Layer 5: 64 $\rightarrow$ 1
Degree=8
166,113
@@
Appendix I Results Details and Visualizations.
@@
In Figure 4, we see that the mean RBA weights for all loss terms eventually converge, indicating mitigation of residual imbalance. In contrast, the GRA weights continue to increase, suggesting persistent gradient imbalance. The steadily growing GRA weights effectively alleviate the gradient stiffness problem, consistent with findings in [63]. The significant magnitude discrepancy between GRA and RBA data justifies using a logarithmic function for GRA weights in loss weighting (Figure 6).
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption

### v2

@@ Appendix D Impact Statement
Appendix E Experiment Setup Details
@@ E.1 Running Time
Model First 5 PDEs Last 3 PDEs 1D- Wave 1D- Reaction 2D NS Cylinder 1D Conv. Diff. Reac. 2D Lid- driven Cavity Hetero- geneous Problem Complex Geometry 3D Point-Cloud PINN 00:21:14 00:09:07 00:15:20 00:15:12 00:06:39 00:23:30 00:01:08 00:49:31 PINNsFormer 00:44:21 00:04:09 00:58:54 02:06:37 14:01:55 00:13:31 QRes 01:41:34 00:02:10 00:24:39 00:25:46 00:13:04 00:20:50 00:01:46 01:32:24 FLS 01:38:01 00:01:29 00:11:51 00:50:26 00:35:48 00:13:38 00:01:08 03:04:41 Cheby1KAN 03:32:10 00:12:08 04:24:59 01:45:37 00:45:20 00:50:45 00:03:21 02:27:27 Cheby2KAN 05:03:18 01:06:54 05:41:42 03:01:40 00:45:15 01:35:40 00:03:27 05:26:42 AC-PKAN 01:13:01 00:15:16 02:21:40 02:01:59 00:51:47 01:13:11 00:01:04 04:54:24 KINN 25:00:20 03:04:19 14:31:42 02:41:49 01:51:44 00:14:07 rKAN 12:44:16 01:21:25 05:19:04 02:06:36 06:21:00 00:16:06 07:53:25 FastKAN 09:35:51 05:51:21 02:04:42 03:22:39 03:37:57 00:17:23 fKAN 08:20:34 00:13:09 03:01:41 01:54:22 00:47:41 00:52:05 00:06:22 04:04:48 FourierKAN 03:33:46 01:21:50 02:48:50 02:08:08 07:40:43 00:18:26 13:36:48
Table 10: Running times ($h$:$m$:$s$) for all eight PDE experiments. First row: Five simpler PDEs; second row: Three more complex cases.
E.2 Complex Function Fitting Experiment Setup Details
@@
E.3 Failure Modes in PINNs Experiment Setup Details
@@
We set the parameters $\lambda_{1}=1$ and $\lambda_{2}=0.01$. Since the system lacks an explicit analytical solution, we utilize the simulated solution provided in Raissi et al. (2019). We focus on the prototypical problem of incompressible flow past a circular cylinder, a scenario known to exhibit rich dynamic behavior and transitions across different regimes of the Reynolds number, defined as $\text{Re}=\frac{u_{\infty}D}{\nu}$. By assuming a dimensionless free-stream velocity $u_{\infty}=1$, a cylinder diameter $D=1$, and a kinematic viscosity $\nu=0.01$, the system exhibits a periodic steady-state behavior characterized by an asymmetric vortex shedding pattern in the cylinder wake, commonly known as the Kármán vortex street. All experimental settings are the same as in Raissi et al. (2019). For more comprehensive details about this problem, please refer to that work.
@@
$X$ (Spatial Coordinates) (5000, 2)(5000,\,2) Contains 5,000 spatial points, each with 2 coordinate values ($x$ and $y$).
$t$ (Time Data) (200, 1)(200,\,1) Contains 200 time steps, each corresponding to a scalar value.
$U$ (Velocity Field) (5000, 2, 200)(5000,\,2,\,200) Contains 5,000 spatial points, 2 velocity components ($u$ and $v$), and 200 time steps. The velocity data of each point is a function of time.
$P$ (Pressure Field) (5000, 200)(5000,\,200) Contains pressure data for 5,000 spatial points and 200 time steps.
@@
$\begin{gathered}\frac{\partial u}{\partial t}+\beta\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}-\rho u(1-u)=0,\qquad\forall x\in[0,2\pi],\;t\in[0,1]\\ \texttt{IC:}\qquad u(x,0)=\exp\left(-\frac{(x-\pi)^{2}}{2(\pi/4)^{2}}\right),\hskip 18.49988pt\texttt{BC:}\qquad u(0,t)=u(2\pi,t)\end{gathered}$ (155)
In this equation, $\beta$ represents the convection coefficient, $\nu$ is the diffusivity, and $\rho$ is the reaction coefficient. Specifically, we set $\beta=1$, $\nu=3$, and $\rho=5$. The reaction term adopts the well-known Fisher’s form $\rho u(1-u)$, as utilized in Krishnapriyan et al. (2021). This formulation captures the combined effects of transport, spreading, and reaction dynamics on the state variable $u$.
@@
$x$ (Spatial Coordinates) (10,201, 1)(10,201,\,1) Represents spatial points uniformly distributed over the domain $0,2\pi]$.
$t$ (Time Data) (10,201, 1)(10,201,\,1) Denotes temporal data spanning the domain $[0,1]$ for solution evolution.
$u$ (Solution Values) (10,201, 1)(10,201,\,1) Contains the computed values of the solution function $u(x,t)$ at corresponding spatial and temporal points.
@@
Out of the total 10,20110,\!201 data points, the dataset was partitioned into training and test sets. The training data includes boundary points (where $x=0$ or $x=2\pi$) and a random sample of 3,0003,\!000 interior points, which were used to compute the loss function during model training. The test data consists of the entire remaining dataset, ensuring comprehensive evaluation of the model’s performance.
@@
$\begin{gathered}\mathbf{u}\cdot\nabla\mathbf{u}+\nabla p-\frac{1}{\text{Re}}\Delta\mathbf{u}=0,\qquad\forall\mathbf{x}\in\Omega,\;t\in[0,T]\\ \nabla\cdot\mathbf{u}=0,\qquad\forall\mathbf{x}\in\Omega,\;t\in[0,T]\\ \texttt{IC:}\qquad\mathbf{u}(\mathbf{x},0)=\mathbf{0}\\ \texttt{BC:}\qquad\mathbf{u}=(4x(1-x),0),\qquad\mathbf{x}\in\Gamma_{1}\\ \mathbf{u}=(0,0),\qquad\mathbf{x}\in\Gamma_{2}\\ p=0,\qquad\mathbf{x}=(0,0)\end{gathered}$ (156)
@@
$X$ (Spatial Coordinates) (10,201, 2)(10,\!201,\,2) Contains 10,201 spatial nodes with $(x,y)$ coordinates spanning the cavity domain.
$U$ (Velocity Field) (10,201, 2)(10,\!201,\,2) Horizontal ($u$) and vertical ($v$) velocity components at $\mathrm{Re}=100$, with no-slip boundary conditions and a moving lid ($y=1$) driving the flow.
$P$ (Pressure Field) (10,201, 1)(10,\!201,\,1) Pressure values normalized with respect to the reference boundary condition.
@@
The training set comprises 3,000 randomly sampled spatial points with associated velocity and pressure values, while the test set evaluates the model on the full dataset of 10,201 nodes. Boundary conditions are explicitly enforced for the moving lid (u=4x(1x),v=0u=4x(1-x),v=0) and stationary walls ($u=v=0$), with the pressure field satisfying the incompressibility constraint.
@@ 2 $\rightarrow$ 512 (Linear, SinAct)
E.4 PDEs in Complex Engineering Environments Setup Details
@@
$\begin{cases}a_{1}\Delta u(\bm{x})=16r^{2}&\text{for }r<r_{0},\\ a_{2}\Delta u(\bm{x})=16r^{2}&\text{for }r\geq r_{0},\\ u(\bm{x})=\frac{r^{4}}{a_{2}}+r_{0}^{4}\left(\frac{1}{a_{1}}-\frac{1}{a_{2}}\right)&\text{on }\partial\Omega,\end{cases}$ (157)
where $r=\|\bm{x}\|_{2}$ is the distance from the origin, $a_{1}=\frac{1}{15}$ and $a_{2}=1$ are the material coefficients, $r_{0}=0.5$ defines the interface between the two materials, and $\partial\Omega$ represents the boundary of the square domain $\Omega=[-1,1]^{2}$. The boundary condition is a pure Dirichlet condition applied uniformly on all four edges of the square.
@@
$\Delta u=0\quad\text{in }\Omega=\Omega_{\text{rec}}\setminus\bigcup_{i=1}^{4}R_{i},$ (158)
@@
$R_{1}$ $=\left\{(x,y):(x-0.3)^{2}+(y-0.3)^{2}\leq 0.1^{2}\right\},$
$R_{2}$ $=\left\{(x,y):(x+0.3)^{2}+(y-0.3)^{2}\leq 0.1^{2}\right\},$
$R_{3}$ $=\left\{(x,y):(x-0.3)^{2}+(y+0.3)^{2}\leq 0.1^{2}\right\},$
$R_{4}$ $=\left\{(x,y):(x+0.3)^{2}+(y+0.3)^{2}\leq 0.1^{2}\right\}.$
@@
$u$ $=0\qquad\text{on }\partial R_{i},\qquad\forall i=1,2,3,4,$ (159)
$u$ $=1\qquad\text{on }\partial\Omega_{\text{rec}}.$ (160)
@@
$\mu(z)\Delta u(\mathbf{x})+k(z)^{2}u(\mathbf{x})=f(\mathbf{x})\quad\text{in }\Omega=[0,1]^{3}\setminus\bigcup_{i=1}^{4}\mathcal{C}_{i},$ (161)
@@
AC-PKAN Linear Embedding: in=2, out=32 ChebyKAN Layers: 4 layers, degree=8 LN Layers: 4 layers, features=64 Output Layer: in=64, out=1 Activation: WaveAct 152,357
QRes Input Layer: in=2, out=128 Hidden Layers: 5 QRes blocks, units=128 Output Layer: in=128, out=1 Activation: Sigmoid 166,017
PINN Layer 1: 2 $\rightarrow$ 256, Activation=Tanh Layer 2: 256 $\rightarrow$ 512, Activation=Tanh Layer 3: 512 $\rightarrow$ 128, Activation=Tanh Layer 4: 128 $\rightarrow$ 1 198,145
PINNsformer d_out=1 d_hidden=128 d_model=8 N=1 heads=2 158,721
FLS Layer 1: 2 $\rightarrow$ 256, Activation=SinAct Layer 2: 256 $\rightarrow$ 256, Activation=Tanh Layer 3: 256 $\rightarrow$ 256, Activation=Tanh Layer 4: 256 $\rightarrow$ 1 132,609
Cheby1KAN Layer 1: 2 $\rightarrow$ 32, Degree=8 Layer 2: 32 $\rightarrow$ 128, Degree=8 Layer 3: 128 $\rightarrow$ 64, Degree=8 Layer 4: 64 $\rightarrow$ 32, Degree=8 Layer 5: 32 $\rightarrow$ 1, Degree=8 129,888
Cheby2KAN Layer 1: 2 $\rightarrow$ 32, Degree=8 Layer 2: 32 $\rightarrow$ 128, Degree=8 Layer 3: 128 $\rightarrow$ 64, Degree=8 Layer 4: 64 $\rightarrow$ 32, Degree=8 Layer 5: 32 $\rightarrow$ 1, Degree=8 129,888
KAN* Layers: 2 $\times$ KANLinear Neurons: 9000 Activation: SiLU 60,000*
rKAN Layer 1: 2 $\rightarrow$ 256, Activation=JacobiRKAN(3) Layer 2: 256 $\rightarrow$ 256, Activation=PadeRKAN[2/6] Layer 3: 256 $\rightarrow$ 256, Activation=JacobiRKAN(6) Layer 4: 256 $\rightarrow$ 128, Activation=PadeRKAN[2/6] Layer 5: 128 $\rightarrow$ 1 165,411
FastKAN* FastKANLayer 1: RBF, SplineLinear 16 $\rightarrow$ 2600, Base Linear 2 $\rightarrow$ 2600 FastKANLayer 2: RBF, SplineLinear 20800 $\rightarrow$ 1, Base Linear 2600 $\rightarrow$ 1 75,418*
fKAN Layer 1: 2 $\rightarrow$ 256, Activation=fJNB(3) Layer 2: 256 $\rightarrow$ 512, Activation=fJNB(6) Layer 3: 512 $\rightarrow$ 512, Activation=fJNB(3) Layer 4: 512 $\rightarrow$ 128, Activation=fJNB(6) Layer 5: 128 $\rightarrow$ 1 132,618
FourierKAN Layer 1: 2 $\rightarrow$ 32 Layer 2: 32 $\rightarrow$ 64 Layer 3: 64 $\rightarrow$ 64 Layer 4: 64 $\rightarrow$ 64 Layer 5: 64 $\rightarrow$ 1 Degree=8 166,113
@@
Appendix F Results Details and Visualizations.
@@
In Figure 4, we see that the mean RBA weights for all loss terms eventually converge, indicating mitigation of residual imbalance. In contrast, the GRA weights continue to increase, suggesting persistent gradient imbalance. The steadily growing GRA weights effectively alleviate the gradient stiffness problem, consistent with findings in (Wang et al., 2021). The significant magnitude discrepancy between GRA and RBA data justifies using a logarithmic function for GRA weights in loss weighting (Figure 6).
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption